The uniform horizontal beam is 6 m long weighs 100 kg force, and is pivoted at a point in the wall. The other end of the beam is supported by a rope which makes 30o with the horizontal. Find (a) the tension in the rope and (b) the vertical and horizontal components of the force exerted by the wall at the pivot


1
Expert's answer
2021-07-19T09:47:36-0400

a)


"T\sin{30}=100(9.8)\\T=1960\ N"

b)


"R_x=1960\cos{30}=1700\ N\\R_y=100(9.8) =980\ N"